OpenStudy (anonymous):

Kim invested a total of $1500 between two bonds. One bond earned 8% per annum and the other bond earned 10% per annum. In one year, Kim earned $132 on her investments. How much did she invest in the bond that earned 10%? someone please give me a detailed explanation ive been stuck for the past 30 minjutes trying to figure out an equation

OpenStudy (anonymous):

....... 2 bonds. therefore x and y is the amount in each. x + y = 1500... 8 percent per annum, so 1.08x is left after a year, 10 percent per annum so 1.10y is left after a year. 132 was earned therefore. x+y = 1500 1.08x + 1.1y = 1500+132. solve the system of equations.

OpenStudy (anonymous):

wait, how did you get 1.08 and 1.10?

OpenStudy (anonymous):

oh. 8% GAIN so you already have 100% because that is the full value, so after a year it is 108%, do you see why?

OpenStudy (anonymous):

Ohhhh yah i think i got it

OpenStudy (anonymous):

ok :)